# SPYGLASS. VII-A. The Demographics and Ages of Small Nearby Young Associations

### v1

@@ 4.2.5 Final Population Demographics
Considering our final stellar membership probabilities in addition to the corrections for binaries and low-mass stars, provided by the following:
@@
where $M_{*}$ is the stellar mass, $c_{M,BIN}$ is the binarity mass correction, and $c_{M,lm}$ is the corrective factor to account for the mass of objects below our detection limit. The total population size $N_{tot}$ and total mass $M_{tot}$ are provided by the sums of $N_{x}$ and $M_{x}$, respectively, across all system primaries and stars without a visible companion. We summarize the results in Table 3. The uncertainties in mass and number are largely systematic, and we adopt a 10% uncertainty following Kerr et al. (2024), which accounts for the typical variation between literature IMF and binarity rates (e.g., Duchêne & Kraus, 2013; Sullivan & Kraus, 2021).
The populations in our sample have masses ranging from 8.9 $M_{\odot}$ and 27 expected members in TOR1B to 117 $M_{\odot}$ and 302 expected members in SCYA-54. Of the 15 top-level populations we provide masses for, eight have $M<33M_{\odot}$, and five (TOR1B, LeoC, AriS, ScuN, and CasE) have $M<21~M_{\odot}$ ($\lesssim 60$ stars), making them smaller than even populations like the 67-member TW Hydrae association (Luhman, 2023). These associations are therefore among the smallest ever characterized (Gagné & Faherty, 2018).
@@ 4.2.6 Virial States
where $M_{cl}$ is the cluster mass, $r_{hm}$ is the half-mass radius, and $\eta$ is a factor related to the mass profile (Portegies Zwart et al., 2010). We use the masses computed in Section 4.2.5 for $M_{cl}$, and set $\eta=5$, which is a value consistent with the broadest association density profiles (Kuhn et al., 2019; Wright et al., 2024). A small $\eta$ favors high $\sigma_{virial}$ and therefore more bound states, allowing us to test whether any of these sparse populations could be bound given the most favorable assumptions.
We compute $r_{hm}$ following SPYGLASS-V, which uses $l$ and $b$ galactic sky coordinates to avoid larger uncertainties in the radial direction, and permit the modeling of small-scale dissolving clusters by fitting bi-variate gaussian to the stellar distribution according to a Kernel Density Estimator. We convert the result from sky coordinates to on-sky distance in pc, and convert from 2D to 3D using the Wolf et al. (2010) conversion factor for a Gaussian. We use the total mass of the population as $M$ in all cases. In populations with a dense core and broad halo, this assumes that all halo stars were at one point in the core. In the sparse populations like those we investigate here, these compact dissolving clusters are often the only plausibly bound structures (e.g. Kerr et al., 2022a, b). We compute $\sigma_{1D}$ using the clipped transverse velocity anomalies of stars within $r_{hm}$ with $P_{fin}>0.5$ that have no evidence of binarity. We subsample that dataset, selecting half of the stars at random, compute a 2-$\sigma$ clipped median, and then set $\sigma_{1D}$ and its uncertainty to the mean and standard deviation of the results, de-convolved with the average uncertainties.
@@ 5 Discussion
Our adopted age measurements use the PARSEC isochronal models, which produce ages consistent with our dynamical and lithium depletion age solutions. However, the uncertainties in the non-isochronal ages, systematic and otherwise, leave the age scale weakly constrained in absolute terms. In our stellar model-independent dynamical ages, which have the largest uncertainties of any method we employ, that uncertainty is partially intrinsic, however the initial size of the population also contributes to the uncertainty, as populations that start out larger produce larger dynamical age biases (see Sec 4.1.2). While small star-forming clouds may be assumed for some small associations we cover, the $6.9\pm 0.5$ Myr old TOR1A has $r_{hm}=20.7$ pc, the second largest in our entire sample. With a velocity dispersion of $\sigma_{1D}=0.49$ pc Myr-1 and a dynamical age of only $4.3\pm 2.9$ Myr, stars in this population would need to be nearly an order of magnitude older than any estimate of its age for stellar dispersal to produce its large size, indicating that its scale is primarily primordial. The existence of such a large population containing only 66 $M_{\odot}$ of stellar mass indicates that a compact initial configuration cannot be relied upon, even in low-mass associations. of near-newborn (<10<10 Myr) young associations to determine whether the initial scale correlates with any dynamical or morphological patterns that persist after formation.
@@
Unlike many recent studies, our dynamical ages lack a clear offset relative to the isochronal ages, which has previously been viewed as a gas dispersal timescale (Miret-Roig et al., 2024). The PARSEC ages are only 1.6 $\pm$ 2.3 Myr older than the dynamical ages on average, and the fact that this value is positive is in large part driven by the older populations in our sample, where PARSEC ages are more consistently older. However, older populations require more precise measurements to achieve the same traceback accuracy at formation, and as such, they are more vulnerable to poor velocity measurements, which produce near-zero times of closest approach and bias the dynamical age young. While we use cuts to mitigate this issue in Section 4.1.2, it is difficult to remove all bad stars, especially in older populations with less complete datasets. This may explain the lower dynamical ages relative to PARSEC among older populations without the need for any systematic offset. The dynamical age bias corrections inspired by Couture et al. (2023) are of a similar order to the offset between the dynamical and adopted ages in SPYGLASS-II, suggesting that the offset computed there arose from a lack of this correction. However, more robust offsets have been presented in more massive populations like Upper Sco (Miret-Roig et al., 2024), suggesting that our low-mass populations were either never globally bound, or became unbound during the star formation process.
@@
Some lithium depletion ages have small uncertainties, but they are highly sensitive to outliers, especially in populations with less complete coverage. This, combined with the uncertain dynamical ages, means that age inaccuracies of order 20% or larger cannot be ruled out. This to better anchor the absolute age scale across a wide range of young stellar associations, such as using the Lithium Depletion Boundary (LDB). The LDB is produced by the rapid onset of Li burning in fully convective stars, where nearly all Li is exhausted almost immediately after the onset of lithium burning in the core. More massive stars reach the necessary temperature for Li burning earlier in their evolution, resulting in high Li abundances below the boundary, and low Li abundances above it. The LDB mass at a given age has little dependence on the initial Li abundance and is consistent across different models, making the LDB a gold standard for association ages (Wood et al., 2023). However, there remain relatively few LDB ages in the solar neighborhood (Binks & Jeffries, 2014; Binks et al., 2021; Couture et al., 2023), due to the need for medium to high resolution spectra of dim stars. It is therefore impractical to compute LDB ages for all associations, however new LDB ages can anchor the isochronal ages, either by re-scaling the isochronal results to align with the LDB age scale, or by comparing the CMD of a new population to that of populations with robust LDB ages. Rottensteiner & Meingast (2024) produced a suite of new empirical isochrones that can be used for this latter purpose, however the ages used in that work are largely based on the same PARSEC isochrones that we consider here. The combination of empirical isochrones like these with reliable LDB ages could be a powerful tool for age measurement.
@@
Most of our populations show an under-abundance of stars more massive than 2-4 M. While a bottom-heavy IMF may exist in these low-mass associations, Gaia surveys of several much more massive populations show a similar pattern. This may indicate that high-mass stars are more difficult to assign to parent associations, likely due to the high binarity rate that often introduces large orbital velocities (Sullivan & Kraus, 2021). We therefore cannot rule out the presence of high-mass stars excluded by this survey. These stars may have a significant effect on boundedness, especially in small populations like TOR1B, which is less massive than many individual stars like Mirzam, a CMaN candidate member (Fossati et al., 2015). A survey for missing high-mass stars will be necessary to confirm our total population masses and determine whether our populations follow a standard Salpeter (1955) IMF in the high-mass regime.
@@
Many questions remain about the origins of these populations, such as whether they truly form in isolation, or out of leftover material from larger star-forming events. We address these issues in our companion paper, SPYGLASS-VII-B, where we trace all associations back to their site at formation, identify co-natal structures that affect the formation environment, and search for small-scale dynamical patterns that connect to features seen in simulations.
@@ 6 Conclusion
We have produced the first demographic overview of 15 young associations identified in SPYGLASS-IV. Using Gaia photometry and astrometry alongside ground-based spectra, we have detected new substructure, and calculated their stellar masses and virial states. We have also produced accurate, self-consistent ages that are supported by isochrones, lithium depletion, and dynamics. Our key findings can be summarized as follows:
@@
    The stellar masses in these associations range from 8.9 $M_{\odot}$ to 117 $M_{\odot}$, and the stellar populations range from 27 to 302 members. This makes these associations among the smallest ever discovered, especially among populations with no clear connection to a larger association.
@@
The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. Funding was received by the Heising-Simons Foundation. RMPK acknowledges the use of computational resources at the Texas Advanced Computing Center (TACC) at the University of Texas at Austin, which was used for the more computationally intensive operations in this project. RMPK acknowledges the staff at the McDonald Observatory, who made the many observations presented in this paper possible. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France. This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France. The original description of the VizieR service was published in Ochsenbein et al. (2000). RMPK thanks Rob Jeffries, who provided helpful advice on the use of the EAGLES module. JSS is grateful for support from NSERC Discovery Grant (RGPIN-2023-04849) and a University of Toronto Connaught New Researcher Award. The authors thank the Sloan Digital Sky Survey V for contributing data to this project. Funding for SDSS-V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, The Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University.

### v3

@@ 4.2.5 Final Population Demographics
Considering our final stellar membership probabilities in addition to the corrections for binaries and low-mass stars, the membership contribution that is implied by the presence of a given star is provided by the following:
@@
where $M_{*}$ is the stellar mass, $c_{M,BIN}$ is the binarity mass correction, and $c_{M,lm}$ is the corrective factor to account for the mass of objects below our detection limit. The total population size $N_{tot}$ and total mass $M_{tot}$ are provided by the sums of $N_{x}$ and $M_{x}$, respectively, across all system primaries and stars without a visible companion. We summarize the results in Table 3. The uncertainties in mass and number are largely systematic, and we adopt a 10% uncertainty following Kerr24, which accounts for the typical variation between literature IMF and binarity rates (e.g., Duchene13; Sullivan21).
The populations in our sample have masses ranging from $M_{\odot}$ and expected members in TOR1B to $M_{\odot}$ and expected members in SCYA-54. Of the 15 top-level populations we provide masses for, have , and (TOR1B, LeoC, ScuN, and CasE) have ($\lesssim 60$ stars), making them smaller than even populations like the 67-member TW Hydrae association (Luhman23). These associations are therefore among the smallest ever characterized (Gagne18BXIII).
@@ 4.2.6 Virial States
where $M_{cl}$ is the cluster mass, $r_{hm}$ is the half-mass radius, and $\eta$ is a factor related to the mass profile (PortegiesZwart10). We use the masses computed in Section 4.2.5 for $M_{cl}$, and set $\eta=5$, which is a value consistent with the broadest association density profiles (Kuhn19; Wright24). A small $\eta$ favors high $\sigma_{virial}$ and therefore more bound states, allowing us to test whether any of these sparse populations could be bound given the most favorable assumptions.
We compute $r_{hm}$ following Kerr24, which uses $l$ and $b$ galactic sky coordinates to avoid larger uncertainties in the radial direction, and permit the modeling of small-scale dissolving clusters by fitting bi-variate gaussian to the stellar distribution according to a Kernel Density Estimator. We convert the result from sky coordinates to on-sky distance in pc, and convert from 2D to 3D using the Wolf10 conversion factor for a Gaussian. We use the total mass of the population as $M$ in all cases. In populations with a dense core and broad halo, this assumes that all halo stars were at one point in the core. In the sparse populations like those we investigate here, these compact dissolving clusters are often the only plausibly bound structures (e.g. Kerr22b; Kerr22a). We compute $\sigma_{1D}$ using the clipped transverse velocity anomalies of stars within $r_{hm}$ with $P_{fin}>0.5$ that have no evidence of binarity. We subsample that dataset, selecting half of the stars at random, compute a 2-$\sigma$ clipped median, and then set $\sigma_{1D}$ and its uncertainty to the mean and standard deviation of the results, de-convolved with the average uncertainties.
@@ 5 Discussion
Our adopted age measurements use the PARSEC isochronal models, which produce ages broadly consistent with our dynamical and lithium depletion age solutions. However, the uncertainties in the non-isochronal ages, systematic and otherwise, leave the age scale weakly constrained in absolute terms. In our stellar model-independent dynamical ages, which have the largest uncertainties of any method we employ, that uncertainty is partially intrinsic, however the initial size of the population also contributes to the uncertainty, as populations that start out larger produce larger dynamical age biases (see Sec 4.1.2). While small star-forming clouds may be assumed for some small associations we cover, the $6.9\pm 0.5$ Myr old TOR1A has pc, the largest in our entire sample. With a velocity dispersion of and a dynamical age of only $4.3\pm 2.9$ Myr, stars in this population would need to be nearly an order of magnitude older than any estimate of its age for stellar dispersal to produce its large size, indicating that its scale is primarily primordial. The existence of such a large population containing only $M_{\odot}$ of stellar mass indicates that a compact initial configuration cannot be relied upon, even in low-mass associations. This motivates future dynamical studies of near-newborn (<10<10 Myr) young associations to determine whether the initial scale correlates with any dynamical or morphological patterns that persist after formation.
@@
Unlike many recent studies, our dynamical ages lack a clear offset relative to the isochronal ages, which has previously been viewed as a gas dispersal timescale (MiretRoig24). The PARSEC ages are only 1.6 $\pm$ 2.3 Myr older than the dynamical ages on average, and the fact that this value is positive is in large part driven by the older populations in our sample, where PARSEC ages are more consistently older. However, older populations require more precise measurements to achieve the same traceback accuracy at formation, and as such, they are more vulnerable to poor velocity measurements, which produce near-zero times of closest approach and bias the dynamical age young. While we use cuts to mitigate this issue in Section 4.1.2, it is difficult to remove all bad stars, especially in older populations with less complete datasets. This may explain the lower dynamical ages relative to PARSEC among older populations without the need for any systematic offset. The dynamical age bias corrections inspired by Couture23 are of a similar order to the offset between the dynamical and adopted ages in Kerr22a, suggesting that the offset computed there arose from a lack of this correction. However, more robust offsets have been presented in more massive populations like Upper Sco (MiretRoig24), suggesting that our low-mass populations were either never globally bound, or became unbound during the star formation process.
@@
Some lithium depletion ages have small uncertainties, but they are highly sensitive to outliers, especially in populations with less complete coverage. This, combined with the uncertain dynamical ages, means that age inaccuracies of order 20% or larger cannot be ruled out. This motivates future efforts to better anchor the absolute age scale across a wide range of young stellar associations, such as using the Lithium Depletion Boundary (LDB). The LDB is produced by the rapid onset of Li burning in fully convective stars, where nearly all Li is exhausted almost immediately after the onset of lithium burning in the core. More massive stars reach the necessary temperature for Li burning earlier in their evolution, resulting in high Li abundances below the boundary, and low Li abundances above it. The LDB mass at a given age has little dependence on the initial Li abundance and is consistent across different models, making the LDB a gold standard for association ages (Wood23). However, there remain relatively few LDB ages in the solar neighborhood (Binks14; Binks21; Couture23), due to the need for medium to high resolution spectra of dim stars. It is therefore impractical to compute LDB ages for all associations, however new LDB ages can anchor the isochronal ages, either by re-scaling the isochronal results to align with the LDB age scale, or by comparing the CMD of a new population to that of populations with robust LDB ages. Rottensteiner24 produced a suite of new empirical isochrones that can be used for this latter purpose, however the ages used in that work are largely based on the same PARSEC isochrones that we consider here. The combination of empirical isochrones like these with reliable LDB ages could be a powerful tool for age measurement.
@@
Most of our populations show an under-abundance of stars more massive than 2-4 M. While a bottom-heavy IMF may exist in these low-mass associations, Gaia surveys of several much more massive populations show a similar pattern. This may indicate that high-mass stars are more difficult to assign to parent associations, likely due to the high binarity rate that often introduces large orbital velocities (Sullivan21). We therefore cannot rule out the presence of high-mass stars excluded by this survey. These stars may have a significant effect on boundedness, especially in small populations like TOR1B, which is less massive than many individual stars like Mirzam, a CMaN candidate member (Fossati15). A survey for missing high-mass stars will be necessary to confirm our total population masses and determine whether our populations follow a standard Salpeter55 IMF in the high-mass regime.
@@
Many questions remain about the origins of these populations, such as whether they truly form in isolation, or out of leftover material from larger star-forming events. We address these issues in our companion paper, Kerr25c, where we trace all associations back to their site at formation, identify co-natal structures that affect the formation environment, and search for small-scale dynamical patterns that connect to features seen in simulations.
@@ 6 Conclusion
We have produced the first demographic overview of 15 young associations identified in Kerr23. Using Gaia photometry and astrometry alongside ground-based spectra, we have detected new substructure, and calculated their stellar masses and virial states. We have also produced accurate, self-consistent ages that are supported by isochrones, lithium depletion, and dynamics. Our key findings can be summarized as follows:
@@
    The stellar masses in these associations range from $M_{\odot}$ to $M_{\odot}$, and the stellar populations range from to members. This makes these associations among the smallest ever discovered, especially among populations with no clear connection to a larger association.
@@
The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. Funding was received by the Heising-Simons Foundation. RMPK acknowledges the use of computational resources at the Texas Advanced Computing Center (TACC) at the University of Texas at Austin, which was used for the more computationally intensive operations in this project. RMPK acknowledges the staff at the McDonald Observatory, who made the many observations presented in this paper possible. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France. This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France. The original description of the VizieR service was published in Vizier00. RMPK thanks Rob Jeffries, who provided helpful advice on the use of the EAGLES module. JSS is grateful for support from NSERC Discovery Grant (RGPIN-2023-04849) and a University of Toronto Connaught New Researcher Award. J.C. acknowledges support from the Agencia Nacional de Investigacioń y Desarrollo (ANID) via Proyecto Fondecyt Regular 1231345, and by ANID BASAL project FB210003. The authors thank the Sloan Digital Sky Survey V for contributing data to this project. Funding for SDSS-V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, The Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University.